OpenStudy (anonymous):

6x-5=8x-50 solve for x cant remember how to factor

OpenStudy (austinl):

You don't need to factor on this problem. You just need to subtract 6x from both sides and add 50 to both sides.

OpenStudy (anonymous):

6x-5=8x-50 -6 -6 -5=2x-50 +50 +50 45=2x 45/2= 22.5

OpenStudy (cggurumanjunath):